OpenStudy (anonymous):

Which equation shows the substitution method being used to solve the system of linear equations? x – y = 7 x = y + 3

OpenStudy (anonymous):

A. x – (y + 3) = 7 B. x = (7 – x) + 3 C. (y + 3) – y = 7 D. x – y = y + 3

OpenStudy (anonymous):

@Catlover5925

OpenStudy (anonymous):

again.... :) as you can tell, im not very good at math. its my worst subject

OpenStudy (midhun.madhu1987):